OpenStudy (anonymous):

Help please??? Find S7 for 3 + 8 + 13 + 18 +… 116 121 126 131 Thanks so much <3

OpenStudy (anonymous):

Are you asking for the 7th term?

OpenStudy (anonymous):

yes, the 7 is supposed to be a subscript not just next to it.

OpenStudy (anonymous):

To find a certain term in the sequence you can use the following \[S_n = \frac{ n }{ 2 }(a+t_n)\] where a is the first term, t_n is the following from sequences \[t_n = a+(n-1) \times d\] where here a = first term, n is the term we are looking for and d = common difference.

OpenStudy (anonymous):

Sorry that should say term in the "series" not sequence

OpenStudy (anonymous):

so the equation would be \[t _{n}=3+(n-1)\times5\] ??

OpenStudy (anonymous):

Looks good, n is the term we are looking for :)

OpenStudy (anonymous):

What am I supposed to do next? I tried working out the problem but I got 33 and that's not an option

OpenStudy (anonymous):

33 is good! But note that we have to plug it into the first formula, at the moment you just found term 7 but not S7 \[t_7 = 33\]\[S_n = \frac{ n }{ 2 }(a+t_n)\] notice we plug in t_7 where it's t_n there are it's t_7 :)

OpenStudy (anonymous):

okay so the equation 7/2(3+33)?

OpenStudy (anonymous):

Looks good

OpenStudy (anonymous):

okay I got C. 126

OpenStudy (anonymous):

sounds good! :)